OpenStudy (anonymous):

HELP! Given the line y=-1/2x+5, which pair of equations represents lines that are parallel, and perpendicular, respectively?

OpenStudy (anonymous):

Here are the possible answers y=-1/2 x-2 and y=-2+1 y=-1/2x+9 and y=2x-1 y=1/2-2 and y =1/2x+7 y=1/2x+4 and y=-1/2x-2

OpenStudy (anonymous):

use this to graph http://www.wolframalpha.com/ the answer second one to find perpendicular find the inverse of slope and perpendicular you need different y-intercept

OpenStudy (anonymous):

*parrellel* you need different y-intercept